OpenStudy (anonymous):

complete the proportion RS/VU=ST/x im am trying to find x. its also the same for this one TU * RT + TV *______

OpenStudy (anonymous):

i know the answer to the second question it is ST

OpenStudy (mani_jha):

What do you've to find in the question?

OpenStudy (anonymous):

i need to find x and the only answers given are a. TV b. TU c. RT d. SV

OpenStudy (mani_jha):

Well, use the similarity concept. AQ is parallel to ST, right? Can you prove that triangles AQR and RST are similar?

OpenStudy (anonymous):

ummmm yes you cna prove that the problem is i dont really understand what its asking exactly

OpenStudy (anonymous):

can*

OpenStudy (mani_jha):

Which points are V and U? I don't see them in the diagram.

Directrix (directrix):

For these triangles to be similar to set up a proportion, 2/5 = 3/6 or 2/6 =3/5 would have to be true. Something is wrong with the diagram or the given lengths. Please check the question and diagram. Thanks.

OpenStudy (anonymous):

i accidentally put the wrong figure up

OpenStudy (anonymous):



OpenStudy (anonymous):

PLEASE HELP ME!!!!!!!!!!!!!!!1

Directrix (directrix):

That means the proportion for which you need x could be RS/VU = ST/UT or RS/UV = ST/VT. So, x could be UT or VT. --------- Looking at the statement you wrote: TU * RT + TV *______ I'm wonderng if you meant to write TU * RT = TV * ____. If "its also the same for this one TU * RT + TV *______" means that the answer to the first part (find x) is the same value that is corrrect for TU * RT = TV * ____, then I would guess ST as the answer. The guess is based on reading information into the stated problem and is just a guess.

OpenStudy (anonymous):

so for the first problem which one would it most likely be?

Directrix (directrix):

I would guess ST as the answer.